OpenStudy (anonymous):

A truck enters a highway driving 60mph. A car enters the highway at the same place 11 minutes later and drives 68mph in the same direction. From the time the car enters the highway, how long will it take the car to pass the truck?

OpenStudy (anonymous):

hello :)

OpenStudy (shane_b):

This one is pretty simple. The first car is traveling at 60mph for 11 minutes...so it travels 11 miles before the next car gets on the highway. The difference in speed between the cars is 68mph-60mph=8mph. All you need to do is determine how long it will take for the second car to travel 11 miles at a speed of 8mph. Can you do that?

OpenStudy (anonymous):

i'm not sure if i've got the right idea, my brain is mush at this point

OpenStudy (shane_b):

We can take it step by step if you like. What part is not clear above?

OpenStudy (anonymous):

how do you find out how long it will take a car to go 11 miles at 8mph

OpenStudy (anonymous):

1 hour and 7.5 minutes?

OpenStudy (anonymous):

the answer must be given in rounded minutes so would 68 minutes be correct in this scenario?

OpenStudy (shane_b):

Not quite :) \[\frac{11 m}{8m/h}=?\]

OpenStudy (anonymous):

that confuses me more

OpenStudy (shane_b):

Ok...before I explain that, how did you get your answer?

OpenStudy (anonymous):

well you go 8 miles in one hour traveling at 8mph, so to go another 3 miles you would, so 60 minutes divided by 8 would be 7.5(minutes) so i just added 60 minutes + 7.5 minutes to get 11miles at 8mph

OpenStudy (anonymous):

i am kind of under time restraint could you please give me the answer and explain after, im afraid ill lose my progress being idle like this (its a form online)

OpenStudy (shane_b):

\[\frac{11 m}{8m/hr}=>11\cancel{m}\frac{1hr}{8\cancel{m}}=1.375hrs\]\[1.375*60=82.5 \space mins=1hr \space 22.5mins=1hr \space 23mins\]